# Mitigating Manipulation and Enhancing Persuasion: A Reflective Multi-Agent Approach for Legal Argument Generation

### v1

@@ 4.2. Scenarios: Arguable, Mismatched, and Non-Arguable Case Triples
”Mismatched” and ”non-arguable” scenarios test ethical AI behavior; abstention is key. An example of case structures for these scenarios is in Table 1.
The generation process for these case triples is parametric, allowing for the specification of several key aspects. These include the total number of case triples to generate, the complexity level (which dictates the number of factors assigned to each case, typically ranging from complexity-1 to complexity+1), and, crucially, the scenario ‘mode‘ (Arguable, Mismatched, or Non-Arguable). For this study, we generated sets of 90 case triples for each of the three scenarios, with a specified complexity level of 5. This means each case within a triple typically contained between 4 and 6 factors.
@@ 5. Results and Analysis
This section presents a detailed empirical analysis of the Reflective Multi-Agent framework’s performance in comparison to the Single Agent, Single Agent with Enhanced Prompting, and Multi-Agent Debate without Reflection baselines. The evaluation is conducted across four distinct LLMs and three specifically designed legal scenarios: ‘arguable,’ ‘mismatched,’ and ‘non-arguable.’ Performance is assessed using three key metrics: Hallucination Accuracy, Factor Utilization Recall, and Successful Abstention Ratio, each addressing critical aspects of argument quality, factual grounding, and ethical AI behavior.
@@ 5.1. Hallucination Accuracy
To assess the models’ ability to avoid generating factually incorrect or unsupported statements (intrinsic hallucination), we measured Hallucination Accuracy. Table 2 presents the Hallucination Accuracy scores for each model and method across the three scenarios. Generally, all configurations achieve high accuracy (often ¿98%) in ‘arguable’ and ‘mismatched’ scenarios. The ‘non-arguable’ scenario, designed to lack a proper basis for argument, posed a more significant challenge to factual fidelity. The Reflective Multi-Agent setup consistently demonstrated superior or near-best performance, notably by substantially improving scores in the ‘non-arguable’ scenario for all tested models. For instance, Reflective Multi-Agent with GPT-4o achieved 94.63% accuracy in this challenging scenario, a marked improvement. This pattern suggests that when no legitimate argument can be formed, methods lacking robust reflective mechanisms are more prone to fabricating factors to fulfill the generation task—a tendency that the Reflective Multi-Agent framework effectively mitigates.
@@
The results in Table 2 confirms that hallucination is more prevalent in the ‘non-arguable’ scenario. The Reflective Multi-Agent method generally shows high hallucination accuracy, with significant improvements in the ‘non-arguable’ scenario. This is critical, indicating the reflective process helps prevent models from fabricating facts when no valid basis exists.
@@ 5.2. Factor Utilization Recall
Effective legal arguments comprehensively incorporate relevant factual information from the provided case materials. Factor Utilization Recall was employed to measure the extent to which each method successfully utilized the ground-truth factors in the ’Arguable’ scenario, where substantive argument generation is expected. The results, detailed in Table 3, indicate that multi-agent configurations (Multi-Agent and Reflective Multi-Agent) generally outperformed single-agent approaches (Single Agent and Single Agent with Enhanced Prompting). The introduction of reflection within the Reflective Multi-Agent framework provided a further discernible improvement over the Multi-Agent baseline. This suggests that both the interactive nature of multi-agent debate and the explicit analytical step of factor usage review contributed to a more thorough incorporation of case factors when arguments were contextually appropriate.
@@
Note: For each model, the highest value among methods for the ‘Arguable’ scenario is bolded. Factor Utilization Recall for ‘Mismatched’ and ‘Non-Arguable’ scenarios is not shown here as successful abstention (detailed in Table 4) is the primary success metric for those scenarios, making recall less informative.
Table 3 indicates that Reflective Multi-Agent achieved high factor utilization in ‘arguable’ scenarios. The removal of ‘mismatched’ and ‘non-arguable’ columns is because successful abstention, as shown in Table 4, is the more relevant metric for those scenarios.
@@ 5.3. Successful Abstention Ratio
A critical characteristic of a reliable and ethically sound AI system is its ability to recognize when an argument is untenable and to appropriately refrain from generating a response. The Successful Abstention Ratio evaluates this capacity, particularly in the ‘mismatched’ and ‘non-arguable’ scenarios where abstention is the desired outcome. As shown in Table 4, the Reflective Multi-Agent configuration demonstrated vastly superior performance in achieving successful abstention compared to all baseline methods.
@@
Table 4 underscores the significant strength of the Reflective Multi-Agent framework in this domain. While Single Agent with Enhanced Prompting and Multi-Agent configurations offered some marginal improvements over the basic Single Agent method, their abstention capabilities remained limited. In contrast, Reflective Multi-Agent consistently achieved significantly higher successful abstention rates across all tested LLMs. This finding strongly suggests that the Factor Analyst agent, with its explicit mandate to assess the groundability of an argument and trigger termination if necessary, played an effective role as a gatekeeper against the generation of inappropriate or unsupportable claims.
@@ 5.4. Synthesis of Findings
The collective results from the evaluation of Hallucination Accuracy, Factor Utilization Recall, and Successful Abstention Ratio revealed a compelling synergistic relationship between the multi-agent architecture and the reflection mechanisms embedded within the Reflective Multi-Agent framework. The Multi-Agent setup, by itself, demonstrated certain benefits over Single Agent approaches, particularly in enhancing factor utilization, likely due to the dynamic of argument and counter-argument construction. However, it was the introduction of explicit, role-based reflection in Reflective Multi-Agent that elevated performance across the most critical aspects related to ethical and reliable argument generation.
Reflective Multi-Agent’s superiority was most pronounced in reducing hallucinations, especially in ‘non-arguable’ contexts where the temptation for models to invent information was high, and in achieving successful abstention when arguments were ungroundable. Importantly, this substantial improvement in successful abstention did not come at the cost of performance in other key areas; the Reflective Multi-Agent framework generally maintained or enhanced hallucination accuracy (Table 2) and factor utilization recall (Table 3) compared to baselines. The Factor Analyst component appeared crucial, acting as a dedicated mechanism for ensuring factual grounding and prompting abstention when warranted. The Argument Polisher, while not directly measured by these quantitative metrics, contributed by refining the output based on the Factor Analyst’s feedback, ensuring that any necessary revisions maintained factual integrity. In essence, while multi-agent debate may contribute to exploring what could be argued, the reflective layer in Reflective Multi-Agent critically controlled how and, crucially, whether to argue. This combination allowed the Reflective Multi-Agent framework to not only generate more comprehensive and factually sound arguments when appropriate but also to reliably abstain when arguments were untenable, thereby addressing key dimensions of the research questions (RQ1 and RQ2) regarding performance and the impact of reflection. Consequently, these improvements in generating well-grounded arguments and appropriately abstaining directly contribute to the development of more ethically persuasive and less manipulative legally compliant intelligent chatbots, a key concern of RQ3.
@@ 6.1. Technical Performance and Ethical Goals
Factor Utilization as Transparent Persuasion: High factor utilization meant arguments were based on provided evidence, making reasoning transparent. Ethical persuasion is evidence-based, aligning with legal principles requiring substantiated arguments. Comprehensive factor utilization was key to persuasive arguments (nadler1983evidence).
Hallucination Accuracy and Abstention as Safeguards Against Manipulation: Generating statements on fabricated factors was informational manipulation (rhode1985ethical). Reflective Multi-Agent’s superior hallucination accuracy, especially in ‘non-arguable’ scenarios, curtailed this. Attempting arguments in ‘mismatched’ or ‘non-arguable’ scenarios misled users. Reflective Multi-Agent’s high abstention rates safeguarded against this, ensuring the system knew when to stop, crucial for responsible AI.
@@ 6.2. Error Analysis: Illustrative Examples of RMA Superiority
Example 1: Hallucination in a ”Arguable” Scenario
@@
Example 2: Failure to Abstain in a ”Non-Arguable” Scenario
@@ 7. Limitations and Future Work
The Reflective Multi-Agent framework promoted ethical persuasion and mitigated manipulation. Reflective Multi-Agent’s evidence-grounded arguments fostered transparency. Enhanced hallucination accuracy and abstention reduced misleading users (tonmoy2024comprehensive). Legal professionals have duties of competence and candor with AI (davis2020future). Reflective Multi-Agent could assist in meeting these. However, AI perpetuating biases in legal data remained a challenge (draper2023potential).
@@ 7.1. Limitations of the Current Research
The current research, while demonstrating the promise of the Reflective Multi-Agent framework, was subject to several limitations that warranted discussion. A primary constraint was the scope of legal factors; the study employed a predefined, discrete set of factors. Real-world legal reasoning, however, is considerably more nuanced and often involved interpreting and weighing factors that were not explicitly enumerated. Consequently, the factor definition and extraction stages, treated here as a pre-process, represented a simplification of a complex interpretative task. Another significant limitation was the depth of reflection embedded in the Reflective Multi-Agent framework. The current implementation involved a single round of analysis and polishing by the Factor Analyst and Argument Polisher, respectively, with at most one revision cycle. While effective, more sophisticated or iterative reflection mechanisms might yield further improvements in argument quality, though this would likely come at an increased computational cost and require careful design to avoid excessive processing times.
@@ 7.2. Directions for Future Research
Building upon the findings and limitations of this work, several promising directions for future research emerged. A key area for advancement was through enhanced factor analysis and knowledge grounding. This could involve developing methods for the system to identify implicit factors not explicitly listed in the input, assess the dynamic relevance of factors based on evolving case narratives, or integrate Retrieval Augmented Generation (RAG) techniques to ground arguments in external legal knowledge bases such as statutes, case law repositories, or scholarly articles (zheng2025reasoning). Such enhancements would move beyond reliance on pre-defined factor sets and allow for more robust and contextually aware reasoning. While the current study demonstrated strong performance using factor-represented inputs, future work should explore the adaptation of the RMA framework to process full-text legal documents directly. This would involve tackling challenges in automated information extraction but could unlock the potential to capture richer contextual nuances from unstructured legal narratives. Another important avenue involved advanced argument polishing and rhetorical control. This could entail fine-tuning the Argument Polisher agent on specialized corpora of legal writing to better capture an appropriate legal style and tone, or developing mechanisms that allowed users to explicitly yet ethically guide the rhetorical strategies employed in the generated arguments (wu2023autogen; sun2024llm), ensuring that persuasiveness did not compromise fairness or accuracy.
@@
Further research should also explore the development of dynamic and adaptive reflection mechanisms. Instead of a fixed number of reflection cycles, the system could be designed to adjust the depth or intensity of reflection based on the initial quality of the generated argument, the complexity of the legal scenario, or even the confidence scores of the agents involved. Designing effective human-in-the-loop collaboration systems was another critical direction, allowing legal professionals to interact with, guide, and iteratively refine the arguments generated by the Reflective Multi-Agent framework. Significant effort must also be dedicated to addressing potential biases, both those that might be present in the input case factors and those that could be introduced or amplified by the LLMs during the generation process. Methodologies for bias detection and mitigation would be crucial for ensuring equitable and just outcomes. Exploring diverse multi-agent configurations, such as introducing a ”Devil’s Advocate” agent to proactively challenge assertions or experimenting with different numbers of agents and interaction protocols (han2024llm), could also lead to more robust and critically examined arguments. Finally, longitudinal studies on the trust and reliance placed on such AI systems by legal professionals were essential to understand their real-world impact, identify potential risks of over-reliance or misuse, and develop guidelines for their responsible deployment in legal practice.
@@ 8. Conclusion
This paper introduced and evaluated a Reflective Multi-Agent framework to improve LLM-generated 3-ply legal arguments, focusing on ethical persuasion and mitigating manipulation. Empirical results showed Reflective Multi-Agent framework’s advantages over baselines. Reflective Multi-Agent framework achieved vastly superior successful abstention, crucial for preventing misleading discourse. It markedly improved hallucination accuracy (especially in non-arguable contexts) and enhanced factor utilization recall. These findings suggest structured reflection (Factor Analyst, Argument Polisher) within a multi-agent debate offered a robust method for guiding LLMs towards more reliable, ethically sound outputs. By analyzing arguments for grounding, factor use, abstention necessity, and polishing for clarity, Reflective Multi-Agent addressed key LLM weaknesses in law. Developing such systems was critical for trustworthy legally compliant intelligent chatbots. While challenges remained, Reflective Multi-Agent’s principles (role specialization, iterative refinement, explicit analysis) provided a promising direction for AI in law, fostering systems that were persuasive, responsible, and ethical.
[…]
@@ Appendix A Detailed Prompts
    • Abstention IS REQUIRED and you MUST output "REQUIRES_ABSTENTION" if EITHER of the following conditions is true for the primary cited precedent:
@@
      Are there any crucial factors from c1 or precedents (c2/c3) that have been overlooked and could significantly strengthen or weaken the argument?

### v2

@@ 4.2. Scenarios: Arguable, Mismatched, and Non-Arguable Case Triples
“Mismatched” and “non-arguable” scenarios test ethical AI behavior; abstention is key. An example of case structures for these scenarios is in Table 1.
The generation process for these case triples is parametric, allowing for the specification of several key aspects. These include the total number of case triples to generate, the complexity level (which dictates the number of factors assigned to each case, typically ranging from complexity-1 to complexity+1), and, crucially, the scenario “mode” (Arguable, Mismatched, or Non-Arguable). For this study, we generated sets of 90 case triples for each of the three scenarios, with a specified complexity level of 5. This means each case within a triple typically contained between 4 and 6 factors.
@@ 5. Results and Analysis
This section presents a detailed empirical analysis of the Reflective Multi-Agent framework’s performance in comparison to the Single Agent, Single Agent with Enhanced Prompting, and Multi-Agent Debate without Reflection baselines. The evaluation is conducted across four distinct LLMs and three specifically designed legal scenarios: “arguable,” “mismatched,” and “non-arguable.” Performance is assessed using three key metrics: hallucination accuracy, factor utilization recall, and successful abstention ratio, each addressing aspects of argument quality, factual grounding, and ethical AI behavior.
@@ 5.1. Hallucination Accuracy
To assess the models’ ability to avoid generating factually incorrect or unsupported statements (intrinsic hallucination), we measured hallucination accuracy. Table 2 presents the hallucination accuracy scores for each model and method across the three scenarios. Generally, all configurations achieve high accuracy (often ¿98%) in “arguable” and “mismatched” scenarios. The “non-arguable” scenario, designed to lack a proper basis for argument, posed a more significant challenge to factual fidelity. The Reflective Multi-Agent setup consistently demonstrated superior or near-best performance, notably by substantially improving scores in the “non-arguable” scenario for all tested models. For instance, Reflective Multi-Agent with GPT-4o achieved 94.63% accuracy in this challenging scenario, a marked improvement. This pattern suggests that when no legitimate argument can be formed, methods lacking robust reflective mechanisms are more prone to fabricating factors to fulfill the generation task—a tendency that the Reflective Multi-Agent framework effectively mitigates.
@@
The results in Table 2 confirms that hallucination is more prevalent in the “non-arguable” scenario. The Reflective Multi-Agent method generally shows high hallucination accuracy, with improvements in the “non-arguable” scenario. This indicates that the reflective process helps prevent models from fabricating facts when no valid basis exists.
@@ 5.2. Factor Utilization Recall
Effective legal arguments comprehensively incorporate relevant factual information from the provided case materials. Factor Utilization Recall was employed to measure the extent to which each method successfully utilized the ground-truth factors in the “Arguable” scenario, where substantive argument generation is expected. The results, detailed in Table 3, indicate that multi-agent configurations (Multi-Agent and Reflective Multi-Agent) generally outperformed single-agent approaches (Single Agent and Single Agent with Enhanced Prompting). The introduction of reflection within the Reflective Multi-Agent framework provided a further discernible improvement over the Multi-Agent baseline. This suggests that both the interactive nature of multi-agent debate and the explicit analytical step of factor usage review contributed to a more thorough incorporation of case factors when arguments were contextually appropriate.
@@
Note: For each model, the highest value among methods for the “Arguable” scenario is bolded. Factor Utilization Recall for “Mismatched” and “Non-Arguable” scenarios is not shown here as successful abstention (detailed in Table 4) is the primary success metric for those scenarios, making recall less informative.
Table 3 indicates that Reflective Multi-Agent achieved high factor utilization in “arguable” scenarios. The removal of “mismatched” and “non-arguable” columns is because successful abstention, as shown in Table 4, is the more relevant metric for those scenarios.
@@ 5.3. Successful Abstention Ratio
One characteristic of a reliable and ethically sound AI system is its ability to recognize when an argument is untenable and to appropriately refrain from generating a response. The Successful Abstention Ratio evaluates this capacity, particularly in the “mismatched” and “non-arguable” scenarios where abstention is the desired outcome. As shown in Table 4, the Reflective Multi-Agent configuration demonstrated vastly superior performance in achieving successful abstention compared to all baseline methods.
@@
Table 4 underscores the strength of the Reflective Multi-Agent framework in this domain. While Single Agent with Enhanced Prompting and Multi-Agent configurations offered some marginal improvements over the basic Single Agent method, their abstention capabilities remained limited. In contrast, Reflective Multi-Agent consistently achieved higher successful abstention rates across all tested LLMs. This finding suggests that the Factor Analyst agent, with its explicit mandate to assess the groundability of an argument and trigger termination if necessary, played an effective role as a gatekeeper against the generation of inappropriate or unsupportable claims.
@@ 5.4. Synthesis of Findings
The results from the evaluation of hallucination accuracy, factor utilization recall, and successful abstention ratio revealed a synergistic relationship between the multi-agent architecture and the reflection mechanisms embedded within the Reflective Multi-Agent framework. The Multi-Agent setup demonstrated benefits over Single Agent approaches, particularly in enhancing factor utilization, due to the dynamic of argument and counter-argument construction. However, it was the introduction of explicit, role-based reflection in Reflective Multi-Agent that elevated performance across the aspects related to ethical and reliable argument generation.
Reflective Multi-Agent’s superiority was most pronounced in reducing hallucinations, especially in “non-arguable” contexts where the temptation for models to invent information was high, and in achieving successful abstention when arguments were ungroundable. Importantly, this improvement in successful abstention did not come at the cost of performance in other key areas; the Reflective Multi-Agent framework generally maintained or enhanced hallucination accuracy (Table 2) and factor utilization recall (Table 3) compared to baselines. The Factor Analyst component acted as a dedicated mechanism for ensuring factual grounding and prompting abstention when warranted. The Argument Polisher, while not directly measured by these quantitative metrics, contributed by refining the output based on the Factor Analyst’s feedback, ensuring that any necessary revisions maintained factual integrity. While multi-agent debate contributed to exploring what could be argued, the reflective layer in Reflective Multi-Agent controlled how and whether to argue. This combination allowed the Reflective Multi-Agent framework to not only generate more comprehensive and factually sound arguments when appropriate but also to reliably abstain when arguments were untenable, thereby addressing key dimensions of the research questions (RQ1 and RQ2) regarding performance and the impact of reflection. Consequently, these improvements in generating well-grounded arguments and appropriately abstaining contribute to the development of more ethically persuasive and less manipulative legally compliant intelligent chatbots, a key concern of RQ3.
@@ 6.1. Technical Performance and Ethical Goals
Factor Utilization as Transparent Persuasion: High factor utilization meant arguments were based on provided evidence, making reasoning transparent. Ethical persuasion is evidence-based, aligning with legal principles requiring substantiated arguments. Comprehensive factor utilization was key to persuasive arguments (nadler1983evidence, 30).
Hallucination Accuracy and Abstention as Safeguards Against Manipulation: Generating statements on fabricated factors was informational manipulation (rhode1985ethical, 36). Reflective Multi-Agent’s superior hallucination accuracy, especially in “non-arguable” scenarios, curtailed this. Attempting arguments in “mismatched” or “non-arguable” scenarios misled users. Reflective Multi-Agent’s high abstention rates safeguarded against this, ensuring the system knew when to stop, crucial for responsible AI.
@@ 6.2. Error Analysis: Illustrative Examples of RMA Superiority
Example 1: Hallucination in an Arguable Scenario
@@
Example 2: Failure to Abstain in a Non-Arguable Scenario
@@ 7. Limitations and Future Work
The Reflective Multi-Agent framework promoted ethical persuasion and mitigated manipulation. Reflective Multi-Agent’s evidence-grounded arguments fostered transparency. Enhanced hallucination accuracy and abstention reduced misleading users (tonmoy2024comprehensive, 42). Legal professionals have duties of competence and candor with AI (davis2020future, 10). Reflective Multi-Agent could assist in meeting these. However, AI perpetuating biases in legal data remained a challenge (draper2023potential, 11).
@@ 7.1. Limitations of the Current Research
The current research, while demonstrating the promise of the Reflective Multi-Agent framework, was subject to several limitations that warranted discussion. A primary constraint was the scope of legal factors; the study employed a predefined, discrete set of factors. Real-world legal reasoning, however, is considerably more nuanced and often involved interpreting and weighing factors that were not explicitly enumerated. Consequently, the factor definition and extraction stages, treated here as a pre-process, represented a simplification of a complex interpretative task. Another limitation was the depth of reflection embedded in the Reflective Multi-Agent framework. The current implementation involved a single round of analysis and polishing by the Factor Analyst and Argument Polisher, respectively, with at most one revision cycle. While effective, more sophisticated or iterative reflection mechanisms might yield further improvements in argument quality, though this would likely come at an increased computational cost and require careful design to avoid excessive processing times.
@@ 7.2. Directions for Future Research
Building upon the findings and limitations of this work, several promising directions for future research emerged. A key area for advancement was through enhanced factor analysis and knowledge grounding. This could involve developing methods for the system to identify implicit factors not explicitly listed in the input, assess the dynamic relevance of factors based on evolving case narratives, or integrate Retrieval Augmented Generation (RAG) techniques to ground arguments in external legal knowledge bases such as statutes, case law repositories, or scholarly articles (zheng2025reasoning, 48). Such enhancements would move beyond reliance on pre-defined factor sets and allow for more robust and contextually aware reasoning. While the current study demonstrated strong performance using factor-represented inputs, future work should explore the adaptation of the RMA framework to process full-text legal documents directly. This would involve tackling challenges in automated information extraction but could unlock the potential to capture richer contextual nuances from unstructured legal narratives. Another important avenue involved advanced argument polishing and rhetorical control. This could entail fine-tuning the Argument Polisher agent on specialized corpora of legal writing to better capture an appropriate legal style and tone, or developing mechanisms that allowed users to explicitly yet ethically guide the rhetorical strategies employed in the generated arguments (wu2023autogen, 44, 41), ensuring that persuasiveness did not compromise fairness or accuracy.
@@
Further research should also explore the development of dynamic and adaptive reflection mechanisms. Instead of a fixed number of reflection cycles, the system could be designed to adjust the depth or intensity of reflection based on the initial quality of the generated argument, the complexity of the legal scenario, or even the confidence scores of the agents involved. Designing effective human-in-the-loop collaboration systems was another direction, allowing legal professionals to interact with, guide, and iteratively refine the arguments generated by the Reflective Multi-Agent framework. Further effort must also be dedicated to addressing potential biases, both those that might be present in the input case factors and those that could be introduced or amplified by the LLMs during the generation process. Methodologies for bias detection and mitigation would be crucial for ensuring equitable and just outcomes. Exploring diverse multi-agent configurations, such as introducing an adversarial agent to proactively challenge assertions or experimenting with different numbers of agents and interaction protocols (han2024llm, 20), could also lead to more robust arguments. Finally, longitudinal studies on the trust and reliance placed on such AI systems by legal professionals were essential to understand their real-world impact, identify potential risks of over-reliance or misuse, and develop guidelines for their responsible deployment in legal practice.
@@ 8. Conclusion
This paper introduced and evaluated a Reflective Multi-Agent framework to improve LLM-generated 3-ply legal arguments, focusing on ethical persuasion and mitigating manipulation. Empirical results showed Reflective Multi-Agent framework’s advantages over baselines. Reflective Multi-Agent framework achieved vastly superior successful abstention, crucial for preventing misleading discourse. It markedly improved hallucination accuracy (especially in non-arguable contexts) and enhanced factor utilization recall. These findings suggest structured reflection (Factor Analyst, Argument Polisher) within a multi-agent debate offered a robust method for guiding LLMs towards more reliable, ethically sound outputs. By analyzing arguments for grounding, factor use, abstention necessity, and polishing for clarity, Reflective Multi-Agent addressed key LLM weaknesses in law. Developing such systems was important for trustworthy legally compliant intelligent chatbots. While challenges remained, Reflective Multi-Agent’s principles (role specialization, iterative refinement, explicit analysis) provided a promising direction for AI in law, fostering systems that were persuasive, responsible, and ethical.
Declaration of Use of AI
The authors declare that no artificial intelligence tools were used in the writing of this manuscript.
@@ Appendix A Detailed Prompts
    • Abstention IS REQUIRED and you MUST output ‘‘REQUIRES_ABSTENTION’’ if EITHER of the following conditions is true for the primary cited precedent:
@@
      Are there any crucial factors from c1 or precedents (c2/c3) that have been overlooked and could strengthen or weaken the argument?